### v1

# ELO-Rated Sequence Rewards: Advancing Reinforcement Learning Models

Abstract
Reinforcement Learning (RL) is highly dependent on the meticulous design of the reward function. However, accurately assigning rewards to each state-action pair in Long-Term RL (LTRL) challenges is formidable. Consequently, RL agents are predominantly trained with expert guidance. Drawing on the principles of ordinal utility theory from economics, we propose a novel reward estimation algorithm: ELO-Rating based RL (ERRL). This approach is distinguished by two main features. Firstly, it leverages expert preferences over trajectories instead of cardinal rewards (utilities) to compute the ELO rating of each trajectory as its reward. Secondly, a new reward redistribution algorithm is introduced to mitigate training volatility in the absence of a fixed anchor reward. Our method demonstrates superior performance over several leading baselines in long-term scenarios (extending up to 5000 steps), where conventional RL algorithms falter. Furthermore, we conduct a thorough analysis of how expert preferences affect the outcomes.
keywords:
Reinforcement Learning, ELO Rating, Ordinal Reward, Reward Decomposition
00footnotetext: This paper has been accepted by DDCLS 2024.
1 Introduction
Deep Reinforcement Learning (RL) has achieved remarkable progress in various fields, including Atari games [7], Go [2, 5], poker [12], and video gaming [4, 6]. However, real-world challenges frequently involve sparse environmental feedback. For example, a Go game may consist of approximately 200 decision states, yet there is only a signal of reward from the last state. This scarcity of feedback requires the careful design of reward functions for RL algorithms. Furthermore, the vast state spaces in many games hinder the ability of agents to learn effectively through self-exploration alone.
This situation often calls for domain experts to design the reward function [29], or use imitation learning for pre-training to guide the agent to learn effective policies [1, 30]. Nevertheless, reward functions designed by human experts face three main issues: Firstly, there may be some errors or hidden flaws in the reward function, leading RL agents to exploit these errors to obtain unreasonable rewards; Secondly, the design by human experts might cap the potential of RL agents, making the agent difficult to surpass the top human experts; Lastly, addressing large-scale problems typically requires a team of experts. These not only incurs high costs but also tends to produce policies that lack distinctive styles or characteristics due to the need for consensus of experts.
To address these issues, we took inspiration from the ordinal utility in economics, which states that people do not accurately calculate the actual value of each decision when making a decision, but obtain a decision by comparing the pros and cons of each strategy [14]. It is reasonable to assume that a RL agent may not require exact numerical rewards for effective learning. The preferences of different trajectories might help train an equally or more powerful RL agent.
Figure 1: Unlike classical RL, preference-based RL has no direct reward. Expert judgment is required when sampling trajectories from the replay buffer, and reward are obtained through expert judgment.
Preference based RL (PBRL) is the embodiment of this idea: the teacher (human expert) provides his/her the preference for the behavior of two agents (or trajectories), and then the agents use this feedback to update their policies respectively (see Figure 1). Compared with designing a reward function for each state-action pair, PBRL can not only greatly reduce the workload of human experts, but it also makes estimated reward less biased and harder to be exploited. Furthermore, it may guide the different styles of the agent through human preferences, so that humans can better understand the reasons why the agent chooses some certain strategies.
However, there are currently three problems with PBRL . First, due to the possibility of experts making mistakes, the update method of PBRL training is extremely unstable; Second, the preference judgment of experts is based on trajectories, but current RL algorithms are more suitable for the dense reward model; Third, PBRL relies solely on an expert’s preference between two trajectories, rather than leveraging the interrelationships among multiple trajectories to enhance learning.
Therefore, this paper proposes a new PBRL baseline algorithm, ELO rating based RL (ERRL). For converting the ordinal reward into cardinal reward, we adopt the ELO ranking system, which is widely used to measure player strength in zero-sum games. Particularly, we consider human experts’ preferences for trajectories as a zero-sum game, in which the preferences of different trajectories can be converted into the ELO rating for model optimization. To address the issue of sparse trajectory feedback, we utilize maximum likelihood estimation, achieving a more stable reward decomposition algorithm compared to least squares estimation. Additionally, we take some engineering optimization methods to increase the algorithm efficiency. In the experiments, the results on the selected Atari games are comparable to the state-of-the-art results in the literature. We further demonstrate the effectiveness of the proposed ERRL algorithm through the ablation experiments.
2 Preliminaries
2.1 Episodic RL with Trajectory Feedback
In classic RL settings, the research focus in Markov Decision Process (MDP) model, MDP general defined as a tuple:$\langle\mathcal{S},\mathcal{A},T,R\rangle$: where $\mathcal{S}$ and $\mathcal{A}$ denote the spaces of environment states and action set of agent. $T(s,a,s^{\prime})=P(s^{\prime}|s,a)$ and $R(s,a)$ denote the probability of state transition to $s^{\prime}$ and the value feedback when action $a$ is taken on state $s$.
The goal of RL is to find a policy $\pi:~{}\mathcal{S}\rightarrow\mathcal{A}$ maximizing long term reward:
$\max G(s_{t},a_{t})=\max{\sum_{k=1}^{T}\gamma^{k}R(s_{t+k},a_{t+k})}$ (1)
Where $\gamma$ is discount factor.
In real-world RL problem, it is difficult for people to get the reward function $R(s,a)$ of the MDP model, we just can get the result of the game. So, in most cases, RL must face the difficulties of trajectory feedback or sparse rewards. A trajectory segment is a sequence of state-action pairs $\sigma\left((s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{k},a_{k})\right)\in(\mathcal{S}\times\mathcal{A})^{k}$, trajectory feedback RL model general defined as a tuple: $\langle\mathcal{S},\mathcal{A},T,\mathcal{R}\rangle$ [17], Unlike the MDP model, the environment only gives an evaluation of the entire trajectory $\mathcal{R}(\sigma)$ at the end of the game.
$\mathcal{R}(\sigma)=\sum_{(s_{i},a_{i})\in\sigma}R(s_{i},a_{i})$ (2)
In order to improve training efficiency, it is necessary to redistribute the entire reward to each state-action pair [15, 17], convert the trajectory feedback RL problem into a standard form of RL problem. Constructing a proxy reward function $\hat{r}$ to estimate reward on state-action pair like MDP model. In practice, we optimize the proxy reward function using the following loss function.
$\mathcal{L}(\hat{r})=\mathbb{E}_{\sigma\in\mathcal{D}}\left[\left(\mathcal{R}(\sigma)-{\sum}_{(s_{i},a_{i})\in\sigma}\hat{r}(s_{i},a_{i})\right)\right]$ (3)
2.2 Ordinal Utility Theory and Cardinal Utility Theory
Cardinal utility theory is a popular concept in microeconomics during the 19th and early 20th centuries [18]. The basic idea is that utility can be measured and summed up. The base unit of utility is called the utility unit. Utility unit can be expressed numerical e.g. $(1,2,3\dots)$, just as the unit of length can be expressed in meters. The mainstream RL reward function $R(s,a)$ is similar to the definition of Cardinality Utility,
Ordinal utility theory was proposed by Hicks and Allen in 1934 [14], and the main point is that: utility as a psychological phenomenon cannot be measured because it is impossible to find a reasonable unit for measurement utility; They believe that what consumers do in the market is not to measure the utility of goods, but to rank among different goods.
Similarly, in some RL problems, it is very difficult to get an accurate Cardinality Reward, but it is easier to obtain the preferences on different trajectories. For example, there are generally 80$\sim$100 decision points in chess, it is difficult to estimate the numerical value of particular states even professional chess players, but the final match result of chess can be judged by even beginner.
2.3 Elo Rating System
The ELO rating system  [21] is an evaluation method created by Arpad Elo to measure the level of players in various zero-sum game competitions. It is a recognized authoritative method for evaluating players’ abilities widely in competitive competitions, such as chess, Go, football, basketball, e-sports and other competitive competitions. Each player has a hidden rating score, which represents the player’s average competitive level. The higher the score, the stronger the player’s ability. However, due to the randomness of the game and the status of the players, there often exists a gap between the on-the-spot scores and the hidden scores. At the ELO rating model, the hidden score needs to add a fluctuation error, to represent the real ability of the on-the-spot performance. The purpose of the ELO rating is to estimate the true abilities (hidden score) of the players through the win-lose relationship as accurately as possible . The final ELO rating model is as follows:
Random variable $\xi$ obey Logistic distribution, denote the gap between on-the-spot scores and the hidden scores:
$P(\xi\leq x)=F(x)=\frac{1}{1+e^{-x}}$ (4)
$R_{A}$ and $R_{B}$ denote the current hidden score of Player A and Player B, let $D=R_{A}-R_{B}$, At this time, the probability of player A defeat player B is:
$P(D)$ $=1/2+F(x)\big{|}_{0}^{\frac{D}{2}}$ (5)
$=\frac{1}{1+e^{\frac{D}{2}}}$
In order to facilitate calculation and to show the ability gap between players intuitively, ELO modifies the variance of the Logistic distribution and set the initial score at 1500.
Assume the win rate of player A is $E(\sigma^{A},\sigma^{B})$:
$E(\sigma^{A},\sigma^{B})=\frac{1}{1+10^{(R_{B}-R_{A})/400}}$ (6)
$\mu$ denote result function, if $\mu(A)=1$ denote player A defeated player B, $\mu(A)=0$ denote player B defeated player A, Then player A’s ELO rating is updated to:
$R_{A}^{\prime}=R_{A}+K(\mu(A)-E(\sigma^{A},\sigma^{B}))$ (7)
2.4 Performance Based RL (PBRL)
PBRL [16] is a typical algorithm using ordinal utility in RL, which divides the entire trajectory into many sub-trajectories with fixed length $k$, and using $\sigma^{A}\succ\sigma^{B}$ indicates that human experts prefer sub-trajectory $\sigma^{A}$ to sub-trajectory $\sigma^{B}$.
If human experts judge $\sigma^{A}\succ\sigma^{B}$, then there is a high probability that $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$:
$R(s^{A}_{1},a^{A}_{1})+R(s^{A}_{2},a^{A}_{2})+\dots+R(s^{A}_{k_{A}},a^{A}_{k_{A}})>R(s^{B}_{1},a^{B}_{1})+R(s^{B}_{2},a^{B}_{2})+\dots+R(s^{B}_{k_{B}},a^{B}_{k_{B}})$
So it is reasonable to assume that under the proxy reward function $\hat{r}$, the probability of an experts making $\sigma^{A}\succ\sigma^{B}$ is:
$\hat{P}(\sigma^{A}\succ\sigma^{B})=\frac{\exp\sum\limits_{i=1}^{i=k_{A}}\hat{r}(s^{A}_{i},a^{A}_{i})}{\exp\sum\limits_{i=1}^{i=k_{A}}\hat{r}(s^{A}_{i},a^{A}_{i})+\exp\sum\limits_{i=1}^{i=k_{B}}\hat{r}(s^{B}_{i},a^{B}_{i})}$ (8)
Under this assumption, store expert’s preference results $\left(\sigma^{A},\sigma^{B}\,\mu\right)$ into the database $\mathcal{D}$, where $\mu$ is the result function, if $\mu(A)=1$ means $\sigma^{A}\succ\sigma^{B}$, $\mu(A)=0$ means $\sigma^{B}\succ\sigma^{A}$, according to the preference, use the cross entropy loss to optimize the proxy reward function $\hat{r}$ and train reinforcement learning through the proxy reward function $\hat{r}$:
$\mathcal{L}(\hat{r})=-\sum_{(\sigma^{A},\sigma^{B}\,\mu)\in\mathcal{D}}\left(\begin{aligned} \mu(A)\log\hat{P}(\sigma^{A}\succ\sigma^{B})+\\ \mu(B)\log\hat{P}(\sigma^{B}\succ\sigma^{A})\end{aligned}\right)$ (9)
3 Methodology
In this section, we introduce our new method, ELO rating based RL. It abandons the numerical reward, sorts the trajectories through ELO rating, and uses a simplified version of the return decomposition method to assign ELO rating to each state-action pair.
3.1 Trail Reward Estimation Via Elo Rating
Given two trajectories $\sigma^{A}$ and $\sigma^{B}$, suppose that in trajectory $\sigma^{A}$ achieves better results than trajectory $\sigma^{B}$ formulating as $\sigma^{A}\succ\sigma^{B}$ (ie. in Chess, player A wins Player B). Denote $\mathcal{R}(\sigma^{X})=\overline{\mathcal{R}(\sigma^{X})}+\epsilon$ as the reward obtained by this trajectory $\sigma^{X}$, where $\overline{\mathcal{R}(\sigma^{X})}$ represents the expected reward of using the trajectory $\sigma^{X}$, $\epsilon$ denote the reward fluctuation. In general, it is very difficult to estimate $\overline{\mathcal{R}(\sigma^{X})}$, either need to consume huge computing resources to fully explore the entire solution space; or hire human experts to design appropriate intermediate reward to guide the agent to explore more efficiently.
Although it is difficult to give the appropriate reward for each trajectory $\mathcal{R}(\sigma^{X})$, it is much simpler to assess the pros and cons of each trajectory. At this time, it is better to use the idea of ordinal utility, using $\mathcal{U}(\sigma)$ to represent the ELO rating of the trajectory $\sigma$, and use $\mathcal{U}(\sigma)$ as the reward of the entire trajectory to guide RL training. When $\sigma^{A}\succ\sigma^{B}$, it can be regarded as player $\sigma^{A}$ defeated player $\sigma^{B}$ ($S_{A}=1$); if the expert cannot judge the pros and cons of the two trajectories, it can be described a draw, $\sigma^{A}\approx\sigma^{B}$ ($\approx$ indicates that the experts believe that the results of the two trajectories are similar), can be regarded as a draw between the two sides ($S_{A}=0.5$) . If $\sigma^{A}\succ\sigma^{B}$ the event $\hat{\mathcal{U}}(\sigma^{A})>\hat{\mathcal{U}}(\sigma^{B})$ will occur with high probability, and the ELO rating $\hat{\mathcal{U}}(\sigma^{A})$ can be updated to:
$E(\sigma^{A},\sigma^{B})=\frac{1}{1+10^{(\hat{\mathcal{U}}(\sigma^{B})-\hat{\mathcal{U}}(\sigma^{A}))/\eta}}$ (10)
$\hat{\mathcal{U}}^{\prime}(\sigma^{A})=\hat{\mathcal{U}}(\sigma^{A})+K(S_{A}-E(\sigma^{A},\sigma^{B}))$ (11)
Where $\eta$ and $K$ are the parameters to adjust the training fluctuation.
3.2 Indirect Redistribution of Long-Term Reward
$\hat{\mathcal{U}}(\sigma)$ is estimating on the entire trajectory, which should be the sum of the ELO rating for all state-action pairs:
$\hat{\mathcal{U}}(\sigma^{X})=\sum_{t=1}^{k_{X}}\hat{U}(s_{t}^{X},a_{t}^{X})$ (12)
In value based RL, the value function itself implies a one-step reward estimate $R(s,a)=G(s,a)-\gamma G(s^{\prime},a^{\prime})$, there is no need to redefine a proxy reward function, We use deep neural network $\hat{G}_{\text{elo}}(s,a)$ to estimate long-term ELO rating on the state-action pair $(s,a)$. The score for the entire trajectory $\sigma^{X}$ is denote as $\hat{\mathcal{U}}(\sigma^{X})$:
$\hat{\mathcal{U}}(\sigma^{X})$ $=\sum_{t=1}^{k_{X}}\hat{U}(s_{t}^{X},a_{t}^{X})$ (13)
$=\sum_{t=1}^{k_{X}}\left[\hat{G}_{\text{elo}}(s^{X}_{t},a^{X}_{t})-\gamma\hat{G}_{\text{elo}}(s^{X}_{t+1},a^{X}_{t+1})\right]$
The long-term reward for each state-action pair is updated as (proved in Appendix A):
$\hat{G}^{\prime}_{\text{elo}}(s^{X}_{t},a^{X}_{t})=\hat{G}_{\text{elo}}(s^{X}_{t},a^{X}_{t})+\frac{K(S_{X}-E_{X})}{k_{X}}$ (14)
3.3 Combining RL and Elo Rating
At the code implementation level, our algorithm is compatible with mainstream RL algorithms. Our algorithm migrates perfectly to any RL algorithm with only a modification to the reward function. In Algorithm 1, we integrate actor-critic RL with Elo rating.
Algorithm 1 AC RL with Elo Rating
1:Initialize $\mathcal{D}\leftarrow\varnothing$
Predictably, the preferences of experts will have a large impact on the results of ordinal-based RL . In addition, the $K$ and $\eta$ parameters of ELO rating will also have a great impact on the results theoretically. We will discuss this problem in detail in Section 4 .
4 Experiments
In this section, we conduct experiments on a suite of Atari benchmark tasks to investigate the empirical performance of our proposed method. We compare our method with several baseline algorithms in the literature, and conduct an ablation study.
4.1 Performance in Atari Benchmark with episodic preference
4.1.1 Setups
To evaluate the behavior of ERRL, we compared its performance with several based algorithms across four Atari tasks. In ordinal based RL, all trajectories record some key events for expert judgment instead of returning a numerical reward. These tasks are long-term feedback problems (up to 5000 steps). We use the open source RL framework RLLib [24] as the baseline, which is an efficient parallel RL framework and adopts the standard preprocessing of Deepmind [7]. All results have been fine-tuned, and detailed settings can be found in Appendix B.
Since our experiment is based on ordinal RL, how to generate the ordinal model is crucial to the learning results. Here we take Pong as an example to illustrate the ordinal production process.
Pong is a game that simulates table tennis (ping-pong). In normal mode, if you shot the ball out of the opponent’s baseline, you will get 1 point; if you don’t catch the opponent’s shot, you will deduct 1 point. The player who reaches 21 points first wins the game. The final trajectory reward is $\mathcal{R}(\sigma)$, its own score minus the opponent’s score. And our evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then $\sigma^{A}\approx\sigma^{B}$, otherwise see the third item.
  3. 3.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
$k$ indicates the length of the trajectory. In order to avoid the trajectory being too long, the game automatically terminates after the step reaches 5000.
The rationale behind this design is twofold: if defeating your opponent directly is not feasible, the strategy shifts towards dragging your opponent into extended play, aiming to hit as many of the opponent’s balls as possible. Conversely, when you can defeat your opponent, the approach is to leverage any means necessary to secure the win. So there are no more restrictions. For more game settings please refer to Appendix B.
4.1.2 Performance in Atari Benchmark
Figure 2: Learning curves on a suite of Atari benchmark tasks with trajectory feedback. The shaded region indicates 95% confidence interval. Due to the trajectory length is indeterminate, we set up an evaluation point every 60 trajectories, about $3\times 10^{4}\sim 3\times 10^{5}$ time-steps.
A comparison of overall performance, as shown in Figure 2, shows that ERRL generally outperforms the baseline algorithms on these several problems. Our method can achieve higher sample efficiency and generate better policies after convergence. Because the returns are sparse, it is difficult for the original PPO to learn useful information. As for the previous reward decomposition function, it is mainly aimed at the decomposition of numerical reward. The reward obtained directly from the conversion of ordinal reward will increase the fluctuation of training. When the agent employs PPO for training with the reward conversion based on ELO Rating, it can learn an effective policy. This is because the ELO rating tends to amplify the performance gap between trajectories. In the long-term RL problems, ELO ratings are more conducive to highlighting the pros and cons of trajectories than the compact real rewards. So that the agent can find useful information.
4.2 Ablation Study
We study the value of $\eta$ in the ERRL algorithm under the condition of a fixed learning rate of $1\times 10^{-4}$. For the value of $K$, we refer to the original design of ELO rating to fix $K=0.04\times\eta$.
As described in Section 3, $\eta$ controls the fluctuation range of the ELO rating, smaller $\eta$ represents greater randomness of the zero-sum game, larger training fluctuations, but training faster; larger $\eta$ represents less randomness, and reduced training fluctuations, but training slower. So we need a suitable $\eta$ to balance training speed and fluctuation.
Figure 3: Learning curves of ERRL in Pong with different choices of hyper-parameter $\eta$. The curves with $\eta=0.01$ correspond to the default implementation.
We perform an ablation study on the hyperparameter $\eta$ that represents the range of fluctuations in the ELO rating. Considering that the time steps of the game will change with the progress of the game, we decided to relating $\eta$ with the time-steps of the trajectory, that is, the $\eta$ actually used in the ELO Rating calculation is $\eta_{\text{elo}}=\eta_{\text{our}}\times\overline{k_{X}}$. For the convenience of writing, we will use $\eta$ to represent $\eta_{\text{our}}$ in the following text if there is no special emphasis.
We evaluate ERRL with a set of choices of hyperparameter $\eta\in\{0.1,0.05,0.01,0.005,0.001\}$. The experimental results shown in Figure 3 show that this hyperparameter has a great impact on RL training. A larger $\eta$ will cause the training fluctuation to be too large and directly collapse, while a smaller $\eta$ will cause the training to be too slow and easily lead to over fitting. An acceptable range of $\eta$ is in interval $[0.05,0.005]$.
From the time-steps of the entire trajectory, regardless of the size of $\eta$, the final time-steps converge to 2100$\sim$2400. It can be seen that under the same preference, if the training can converge, the choice of $\eta$ does not change the final strategies, and we will discuss the relationship between strategy selection and expert preference in Appendix C.
5 Related Work
RL without Environmental Reward. While RL typically relies on environmental reward, several studies have explored RL without direct environmental rewards. This research can be broadly categorized into two approaches. The first approach focuses on capturing essential information within trajectories and using these differences to define learning rewards Eysenbach et al. [28] proposed a pre-training method to train the discriminator to identify which agent generated the trajectory. the agent always hopes that the discriminator can easily identify itself, so the agent always tends to find a unique state. Hu et al. [26] uses multidimensional features to determine the Pareto frontier. Only the agents at the Pareto frontier will perform the next training. This approach enables game designers to create a variety of agents with distinct styles, but this type of method is generally only suitable for increasing the diversity of agents. Another idea is to use human embeddings in RL, such as PBRL [16] introduced in Section 2.4. In addition, Kimin et al. [13] proposed the method of PEBBLE, which has demonstrated exceptional outcomes. .
Sparse Rewards RL. Sparse rewards significantly hinder the application of RL [11, 8], with numerous studies addressing this challenge. A straightforward solution is the integration of human experience [1, 4]. As long as the human benefit is designed perfectly enough, this approach can lead to successful outcomes. However, relying heavily on human expertise can be expensive and may inadvertently limit the capabilities of the RL agent [1, 4, 5]. Another approach is reward decomposition, which decomposes the total reward into individual state-action pairs [15, 20]. usually this method requires reward shaping to achieve good results. Marcin et al. [10] also achieves good results by reusing historical information, but this only works for tasks with simple and clear states.
6 Conclusion and Future Work
We propose a new preference-based RL algorithm ERRL, which can achieve good performance in the 5000-step long-tern trajectory feedback RL problem. ERRL abandons the classical numerical reward, and estimates the ordinal reward of each trajectory through the expert’s preference and the ELO rating method. In addition, a new reward decomposition method is adopted to reduce the fluctuation of training in ERRL setting. Experiments show that ERRL outperforms previous methods in solution quality. In future work, we will explore the theoretical basis for combining ELO rating with RL, conduct more detailed experiments, discuss the effectiveness of our algorithm when it is difficult for experts to make judgments, and explore the feasibility of ELO rating in the remaining areas. Such as reward shaping  [27], games with hard-to-determine payoffs (ie. Texas hold’em and Hanabi) [25], or research on the diversity of RL agents [26, 28].
Appendix A
$\hat{\mathcal{U}}(\sigma^{A})$ is the ELO rating estimate of the entire trajectory $\sigma^{A}$, in trajectory $\sigma^{A}$, let the long-term ELO rating $\hat{U}(s_{t}^{X},a_{t}^{X})$ for each state-action pair update as $\delta_{t}^{X}$,
$\hat{U}^{\prime}(s_{t}^{A},a_{t}^{A})=\hat{U}(s_{t}^{A},a_{t}^{A})+\delta_{t}^{A}$ (15)
In theory, any update can be regarded as an effective update range as long as it is randomly sampled so that the $\delta_{t}^{X}$ distribution satisfies the $K(S_{X}-E_{X})=\delta_{t}^{X}$ constraint. But this obviously leads to too large variance, which is not suitable for iterative calculation. So we might as well consider $\delta_{t}^{A}$ as the assumption of the ELO rating on the player’s ability to fluctuate, and assume that the estimation error for each state-action pair is independent and uniformly obeys the Logistic distribution:
$\delta_{t}^{A}\sim\text{Logistic}(x)=\frac{e^{-x}}{(1+e^{-x})^{2}}$ (16)
According to maximum likelihood estimation, the above problem can be transformed into an optimization problem. The current formula  17 takes the maximum value of $x_{t}$ as the most likely update range:
$\begin{split}\max\prod_{t=1}^{k_{A}}\frac{e^{-\delta_{t}^{A}}}{(1+e^{-\delta_{t}^{A}})^{2}}\\ \text{s.t.}~{}~{}~{}~{}(S_{A}-E_{A})=\sum_{t=1}^{k_{A}}\delta_{t}^{A}\end{split}$ (17)
Taking the logarithm of the original formula 17, there are:
$\max-\sum_{t=1}^{k_{A}}\left(\delta_{t}^{A}+2\ln(1+e^{-\delta_{t}^{A}})\right)$ (18)
Bringing into $\sum_{t=1}^{k_{A}}=(S_{A}-E_{A})$
$~{}~{}~{}\sum_{t=1}^{k_{A}}\left(\ln(1+e^{-\delta_{t}^{A}})\right)$ (19)
$=\sum_{t=1}^{k_{A}}\left(\frac{\ln(1+e^{\delta_{t}^{A}})}{e^{\delta_{t}^{A}}}\right)$
Finally:
$\begin{split}\min\sum_{t=1}^{k_{A}}\ln(1+e^{\delta^{A}_{t}})\\ \text{s.t.}~{}~{}~{}~{}K(S_{A}-E_{A})=\sum_{t=1}^{k_{A}}\delta_{t}^{A}\end{split}$ (20)
$\begin{split}\min\sum_{t=1}^{k_{A}}\ln(1+e^{\delta^{A}_{t}})\\ \text{s.t.}~{}~{}~{}~{}K(S_{A}-E_{A})=\sum_{t=1}^{k_{A}}\delta_{t}^{A}\end{split}$ (21)
It is easy to know that when $\delta_{t}^{A}=\frac{K(S_{A}-E_{A})}{k_{A}}$, the cost function obtains the minimum value, so the new model is updated as:
$\hat{G}_{\text{elo}}^{\prime}(s_{t}^{A},a_{t}^{A})=\gamma\hat{G}_{\text{elo}}(s_{t+1}^{A},a_{t+1}^{A})+U(s_{t}^{A},a_{t}^{A})+\delta_{t}^{A}\$ (22)
Bringing into $U(s_{t}^{A},a_{t}^{A})=\hat{G}_{\text{elo}}(s_{t}^{A},a_{t}^{A})-\gamma\hat{G}_{\text{elo}}(s_{t+1}^{A},a_{t+1}^{A})$,
$\hat{G}_{\text{elo}}^{\prime}(s_{t}^{A},a_{t}^{A})=\hat{G}_{\text{elo}}(s_{t}^{A},a_{t}^{A})+\delta_{t}^{A}$ (23)
Appendix B
Hyper-Parameter Configuration For Atari
In terms of hardware configuration, the CPU is AMD thread-rippers 3990WX with 32 cores and 64 threads, the graphics card is Nvidia RTX 3080Ti, and the memory is 128GB . During training, 60 workers are used to collect data at the same time to improve the experimental efficiency. The training method of agent is PPO . We refer to the default settings of PPO [24, 22], the main hyperparameter settings are shown in the figure 1.
Table 1: Hyper-Parameter
Hyper-Parameter Default Configuration
discount factor $\gamma$ 0.99
activation ReLU
stacked frames in agent observation 4
state mode memory
actor agent num 60
network architecture PPO
optimizer Adam
learning rate $1\cdot 10^{-4}$
gradient steps per environment step 30000$\sim$300000
ELO $\eta$ 0.01
ELO $K$ $0.04\times\eta$
Our open source main experimental code and details can be found in https://github.com/Zealoter/ERRL, as our experiments are still ongoing, we will gradually improve our code and experimental details.
Expert Preferences
Pong
In Pong game, you control the right paddle, you compete against the left paddle controlled by the computer. Unlike reality ping-pong, the upper and lower sides of the table have boundaries, regardless of the ball going out of bounds. Evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then $\sigma^{A}\approx\sigma^{B}$, otherwise see the third item.
  3. 3.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
SpaceInvaders
In this game, your goal is to fire laser cannons at space invaders to destroy them before they reach Earth. The game is over when you lose all your lives under enemy’s fire, or when they reach Earth.Since the aliens will gradually approach the earth unless they are eliminated. Obviously, the score is not as important as the time in this game. This is because the agent may give up killing the front row in order to kill the high-scoring aliens in the back row. So the evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
  1. 1.
    If $K_{A}>K_{B}$ or $K_{A}<K_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
BeamRider
In this game, you control a space-ship that travels forward at a constant speed. You can only steer it sideways between discrete positions. However, at the end of the level only torpedoes can defeat the boss, and the number of torpedoes is limited, so the preferences of this game are difficult to design. From the player’s point of view it’s always beneficial to live longer, so we combined time-steps length and score to design BeamRider’s preferences. The evaluation criteria for trajectory $\sigma^{A}$ and trajectory $\sigma^{B}$ are:
  1. 1.
    If $\mathcal{R}(\sigma^{A})+K_{A}>\mathcal{R}(\sigma^{B})+K_{B}$ or $\mathcal{R}(\sigma^{A})+K_{A}<\mathcal{R}(\sigma^{B})+K_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Appendix C
Here we discuss several settings of expert preference using Pong and explore the impact of different expert preferences on training.
Preference 1: normal mode
This mode is the same as the previous.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then $\sigma^{A}\approx\sigma^{B}$, otherwise see the third item.
  3. 3.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Preference 2: reward mode
Figure 4: It can be seen that the preference has a great influence on ERRL, and changing the preference may change the final policy of the agent.
In this mode, preference comes only from reward.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Preference 3: ball control mode
In this mode, trajectories with more time-steps are always better in the same reward. That’s why it’s called ball control mode.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Preference 4: aggressive mode
We want the agent to beat the opponent as quickly as possible, so when agent is able to beat the opponent, less time-steps is the better.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then see the third item, if $\mathcal{R}(\sigma^{A})<0$ and $\mathcal{R}(\sigma^{B})<0$ then see the fourth item, $\sigma^{A}\approx\sigma^{B}$.
  3. 3.
    If $k_{A}<k_{B}$ or $k_{A}>k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
  4. 4.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
Preference 5: sudden death mode
In this mode, if $\mathcal{R}(\sigma^{X})\leq-5$, game over directly, and the rest is the same as preference 1.
  1. 1.
    If $\mathcal{R}(\sigma^{A})>\mathcal{R}(\sigma^{B})$ or $\mathcal{R}(\sigma^{A})<\mathcal{R}(\sigma^{B})$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise see the second item.
  2. 2.
    If $\mathcal{R}(\sigma^{A})>0$ and $\mathcal{R}(\sigma^{B})>0$ then $\sigma^{A}\approx\sigma^{B}$, otherwise see the third item.
  3. 3.
    If $k_{A}>k_{B}$ or $k_{A}<k_{B}$ then $\sigma^{A}\succ\sigma^{B}$ or $\sigma^{A}\prec\sigma^{B}$, otherwise the final result is $\sigma^{A}\approx\sigma^{B}$.
It can be seen that inappropriate preferences may fail to train useful agents. In the case of ball control mode, the agent gets a policy that can play more time-steps in the game. However, the rest of the models did not reach satisfactory conclusions, so we will explore the relationship between expert preference and final strategy in future research.